# CulturePark: Boosting Cross-cultural Understanding in Large Language Models

### v1

@@ 4.1 Evaluation on content moderation tasks
Setup. Content moderation is crucial to maintaining the integrity and safety of online platforms in different cultures. What is acceptable in one culture could be offensive or inappropriate in another. However, few methods focused on content moderation for different cultures. For this experiment, we evaluated the effectiveness of our cultural-specific models for $8$ different cultures: Arabic, Bengali, Chinese, German, Korean, Portuguese, Spanish, and Turkish culture. These cultures have their unique characters, involving a large number of people in the world.
To be specific, we evaluate on $7$ content moderation tasks for $8$ different cultures to detect the following content: hate speech, offensive language, spam speech, abusive speech, bias speech, threat speech, and stance of speech in zero-shot evaluation, whose metric is average F1 score. The details on the datasets can be found in Appendix B. In total, our test set contains $48,895$ samples. We compare our models with seven baselines: GPT-3.5-turbo [39], GPT-4 [40], Gemini-pro [41], SeaLLM [18], TaiwanLLM [21], CultureLLM [22], and CultureBank [35]. CultureLLM is a series of culturally specific LLMs using semantic data augmentation. SeaLLM specifically focuses on Southeast Asian (SEA) culture, which is adopted for Chinese and Korean cultures. TaiwanLLM focuses on traditional Chinese culture. CultureBank collects data from social media and we compare on the Arabic and Korean culture by fine-tuning GPT-3.5-turbo on its dataset.
@@
Main Results. We analyzed the results from the culture and task sides in Figure 3. The most interesting observation is that our models outperformed GPT-4 on $5$ cultures and approached GPT-4 on the remaining $3$ cultures, though the data for fine-tuning is generated by GPT-3.5-turbo, which is much worse than GPT-4. We also generated cultural data via GPT-4 and fine-tuned other $8$ cultural-specific models for comparison, denoted as “Ours-gpt4” in Figure 3. The performance of those models is better than “Ours” (GPT-3.5-turbo version) but not so much. For other baselines, our models outperform them in most cases. Table 1 shows that our models achieved better performance than those costly LLMs which require pre-training and fine-tuning.
Ablation study. Ablation study on $4$ cultures is in Table 2, where “Generate” means just extracting opinions from the dialogue, “Verify” represents factually verifying the extracted opinions, and “Diversify” means removing redundant data. Results show that each part of CulturePark is effective.
@@ 4.2 Evaluation on cultural alignment via Hofstede’s Cultural Dimensions Theory
Results. We compared our models (powered by GPT-3.5-Turbo) with GPT-3.5-turbo and GPT-4. As shown in Figure 4, our models outperform GPT-3.5-turbo and GPT-4 by a large margin, indicating their excellent cultural alignment and cultural understanding abilities. Note that VSM is widely adopted as datasets for value and culture alignment, the results imply that our approach for data collection is effective, thus it could be further used for value alignment research.
@@ 4.3 Evaluation in situated learning for cultural education
Situated learning suggests that learning is best understood and facilitated when it occurs within the context [42, 43]. Motivated by situated learning, we leveraged CulturePark for cultural education where our fine-tuned models serve as foreigners to talk to people about cultural problems, which can create a situation for cross-cultural communication and learning cultural-specific knowledge. For example, a person who wants to learn about Arabic culture can communicate with our Arabic model.
Setup and study process. We hired $24$ participants, each of whom was given an outline for cultural learning and asked to chat with models based on the outline. They can ask any related questions and express their opinions to models. Afterwards, participants took a cultural understanding examination from VSM 2013 [1, 26] which they had never come into contact with. We then computed the Euclidean distance between the ground truth and their answers from six cultural dimensions (rf. Section 4.2). For comparison, $12$ participants learned with our models and the other $12$ learned with GPT-4 to study $6$ cultures: Arabic, Bengali, German, Korean, Portuguese and Spanish culture. Each culture was learned by four participants, where two of them learn with our models, and the others learn with GPT-4. The detailed information of participants can be found in table 9. During the study process, first, each participant was given an outline for cultural learning written by cultural experts. The outline (Section C.2), serves as the guideline for efficient learning. Then, we asked the participants to freely communicate with the models to learn about specific cultures. After the examination, we asked the participants to give a score of 1-5 to indicate their satisfaction with the learning process. Throughout this study, we aim to answer two questions: 1) What is the learning performance of the participants with our models and GPT-4? 2) How are their learning experience?
@@
Results. Table 3 shows the averaged results from different participants. We have the following findings. First, participants learning with our models achieved better performance in cultural examination than those with GPT-4 in all cultures. This indicates that our fine-tuned models have a better cultural understanding than GPT-4. Second, participants are more satisfied with communicating with our models than GPT-4. Furthermore, many participants expressed that the responses from GPT-4 are vague. Even though we have prompted GPT-4 to be like a person from a specific culture, it always responds with neutral words that have no clear opinions or ideas. Instead, our models can provide straightforward opinions.
@@ 5.1 Why CulturePark benefits fine-tuning?
Cross-cultural communication vs. direct generation from GPT-4 / GPT-3.5. We compared the results of fine-tuning using data directly generated by GPT models (i.e., no communication). We generated such data by prompting GPT-4 or GPT-3.5-turbo as: “Question: {input} Answer: {output} Please list 10 reasons to support the answer and number them”. Then, these data are used to fine-tune GPT-3.5-turbo. Figure 5(a) shows the performance on content moderation tasks in Chinese, Korean and Turkish cultures. We see that data directly generated from GPT-4 is better than that from GPT-3.5, while our GPT-3.5-based models can outperform both of them.
Diversity of the generated data. We also analyzed the diversity gain [44] of the generated dataset to evaluate the data quality. We compared with CultureLLM [22] and presented the results in Figure 5(b). It indicates that CulturePark can generate more diverse and high-quality data.
(a) Benefits of communication.
(b) Diversity gain compared with CultureLLM [22].
(c) Influence of cultural background and gender.
(d) Performance on BBH.
Figure 5: More discussions on CulturePark.
@@ 5.2 Exploring agents’ cultural background and gender
To explore the influence of agent’s cultural background and gender, we conducted three types of multi-agent communications in Arabic culture: “In-cultural+Different gender”, “Cross-cultural+Same gender”, and “Cross-cultural+Different gender”. [^6] For a fair comparison, the number of seed data for this part is $50$ from WVS [28] following CultureLLM [22]. For each setting, we fine-tuned three different models, whose training data is $500$, $750$, and $1000$, representative. We evaluated models’ performance on content moderation tasks and presented the results in Figure 5(c). “Cross-cultural+Different gender” exhibits the best performance and ability to generate more high-quality data. This indicates the necessity of bringing more diversity in data generation, as conducted in CulturePark.
@@
A potential dilemma arises when fine-tuning LLMs for specific tasks, as it may result in catastrophic forgetting of its original capabilities. This section explores the extent of forgetting exhibited by CulturePark on BIG-Bench-Hard (BBH) [45], which comprises 21 tasks that assess semantic understanding and logical reasoning. For cost efficiency, we sampled $100$ samples from each BBH task. We evaluated our models against the baseline model, GPT-3.5-turbo. The results in Figure 5(d) indicate that CulturePark generally maintains or even enhances performance on most benchmarks, including the 21 tasks in BBH. This improvement suggests potential latent relationships between cultural data and general benchmarks, implying that fine-tuning on cultural data could enhance general reasoning abilities.
@@ 5.4 Open-source fine-tuning with Llama2-70b
To verify the generalization ability of our framework, we leveraged the generated data to fine-tune cultural specific Llama-2-70b models and evaluate on content moderation tasks. As shown in Figure 6, our models outperform Llama-2-70b in all $8$ cultures, especially in German, Chinese, Bengali, and Portuguese cultures, which cover both low- and high-resource cultures. Furthermore, our models are also excellent in all $7$ tasks. [^7] The detailed hyperparameter setting is shown in Table 6. We only conduct this pilot study to show that Llama2 also works in CulturePark. The results might be further improved with more GPU resources. This verifies the generalization of CulturePark as an effective data collection platform.
@@ 6 Conclusions, Societal Impact, and Limitations
This paper introduced CulturePark, an LLM-powered multi-agent framework for cultural data collection through multi-agent communication. CulturePark can generate high-quality and diverse cross-cultural dialogue, which can be used to fine-tune cultural specific LLMs. We evaluated CulturePark across three downstream tasks: content moderation, cultural alignment, and cultural education, indicating great improvement over GPT-4.
@@
Our work has the following limitations. 1) More experiments can be done by replacing GPT-3.5-Turbo in CulturePark to discover more results. 2) Our fine-tuned models are mostly for high-resource cultures. The reason is that the dataset and benchmark on low-resource cultures are rare, and we can not find enough data for fine-tuning and evaluation. 3) More efficient fine-tuning techniques can be studied to support the fine-tuning of culturally specific LLMs.
@@
The human study was conducted following local laws and regulations, and the evaluation process was controlled to ensure that no irresponsible content was generated. Authors respect all cultures studied in the world. Results in the paper may change due to the change of OpenAI API and their model versions.
Appendix A Details on the cross-cultural dialogue
A.1 Examples on seed data
@@ A.2 The dialogue dataset
Then, we randomly sampled $750$ dialogues for each culture and evaluated the communication using prompts in Appendix D. As shown in Table 4, on average, the average ratio of statements that express cross-cultural understanding is $80.80$%. The analysis also verifies the effectiveness of CulturePark in extending topics and cross-cultural understanding.
A.3 Details on data refinement
Algorithm 1 shows the algorithm pipeline of data refinement.
Input: a dialogue record $record$, targeted culture $culture$, seed data $s$
Output: opinions of targeted culture $Final\_opinions$
$Embeddings\leftarrow[]$;
$opinions\leftarrow[]$;
for $r$ in Records do
       if $r$ is from $culture$ then
             extract $opinion$ from $r$ via GPT-4;
             judge the $relationship$ of $s$ and $opinion$ from Entail, Contradict and Irrelevant;
             if $relationship$ == Entail then
                   append $opinion$ into $opinions$;
                   $embedding$ = getEmbedding($opinion$);
                   append $embedding$ into $Embeddings$;
            else if $relationship$ == Contradict then
                   rewrite the $opinion$ and check the $relationship$ again;
                   if $relationship$ == Entail then
                         append $opinion$ into $opinions$;
                         $embedding$ = getEmbedding($opinion$);
                         append $embedding$ into $Embeddings$;
$Labels$ = clustering($Embeddings$, min(10, len($Embeddings$)));
$Final\_opinions\leftarrow[]$;
for $arr$ in $Labels$ do
       $final\_opinion$ = randomly select one from $arr$;
       append $final\_opinion$ into $Final\_opinions$;
return $Final\_opinions$;
Algorithm 1 Cultural data refinement
@@ A.3 Details on data refinement
Appendix B Details on the test sets
@@
Table 5: A brief introduction of the $8$ evaluation tasks and $41$ datasets. We list both the name and the size of test sets. For instance, “OSACT5(2541) [46]” denotes that there are 2541 test samples in the dataset OSACT5.
@@
Arabic (CulturePark-Ar) Middle East Offensive language detection: OSACT5(2541) [46]. Hate detection: Multi-Platform(675) [47], OSACT5(2541) [46], and OSACT5_finegrained(2541) [46]. Vulgar detection: Multi-Platform(675) [47] 8,973
@@
Bangli (CulturePark-Bn) Bangladesh Offensive language detection: TRAC2020 Task1(1000) [48], TRAC2020 Task2(1000) [48]. Hate detection: Hate Speech(1000) [49]. Threat detection: BACD(1000) [50]. Bias detection: BACD(1000) [50]. 5,000
@@
Chinese (CulturePark-Zh) China Spam detection: CCS(1000) [51]. Bias detection: CDial-Bias(1000) [52]. 2,000
@@
German (CulturePark-De) Germany and parts of Europe Offensive language detection: GermEval2018(3531) [53]. Hate detection: IWG_1(469) [54], IWG_2(469) [54], HASOC2020(850) [55], and multilingual-hatecheck(1000) [56]. 6,319
@@
Korean (CulturePark-Ko) South Korea Hate detection: hateSpeech(1000) [57], and HateSpeech2(1000) [58]. Abusive detection: AbuseEval(1000) [59], CADD(1000) [60], and Waseem(1000) [61]. 5,000
@@
Portuguese (CulturePark-Pt) Brazil and parts of Latin America Offensive language detection: OffComBR(1250) [62], and HateBR(1000) [63]. Bias detection: ToLD-Br-homophobia(1000) [64], and ToLD-Br-misogyny(1000) [64]. Abusive detection: ToLD-Br-insult(1000) [64]. 5,250
@@
Spanish (CulturePark-Es) Argentina, Mexico, and parts of Latin America Offensive language detection: AMI(1000) [65], MEX-A3T(1000) [66]. Hate detection: HatEval 2019(1000) [67]. Bias detection: DETOXIS_stereotype(1000) [68], and DETOXIS_improper(1000) [68]. Abusive detection: DETOXIS_abusive(1000) [68], DETOXIS_mockery(1000) [68]. Aggressiveness detection: DETOXIS_aggressiveness(1000) [68]. Stance detection: DETOXIS_stance(1000) [68]. 9,000
@@
Turkish (CulturePark-Tr) Turkey Offensive language detection: SemEval-2020(3528) [69], and offenseKaggle_2(1000) [70]. Abusive detection: ATC(1000) [71]. Spam detection: Turkish Spam(825) [72]. Fine-grained offensive detection: offenseCorpus(1000) [73]. 7,353
B.1 Arabic

### v2

@@ 4.1 Evaluation on content moderation tasks
Setup.
Content moderation is crucial to maintaining the integrity and safety of online platforms in different cultures. What is acceptable in one culture could be offensive or inappropriate in another. However, few methods focused on content moderation for different cultures. For this experiment, we evaluated the effectiveness of our cultural-specific models for $8$ different cultures: Arabic, Bengali, Chinese, German, Korean, Portuguese, Spanish, and Turkish culture. These cultures have their unique characters, involving a large number of people in the world.
We evaluate on $7$ content moderation tasks for $8$ different cultures to detect the following content: hate speech, offensive language, spam speech, abusive speech, bias speech, threat speech, and stance of speech in zero-shot evaluation, whose metric is average F1 score. The details on the datasets can be found in Appendix C. In total, our test set contains $48,895$ samples. We compare our models with seven baselines: GPT-3.5-turbo (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), Gemini-pro (Google, 2023), SeaLLM (Nguyen et al., 2023b), TaiwanLLM (Lin and Chen, 2023), Synatra-7B-v0.3-dpo (maywell, 2024), EEVE-Korean-10.8B-v1.0 (yanolja, 2024), CultureLLM (Li et al., 2024), and CultureBank (Shi et al., 2024). CultureLLM is a series of culture-specific LLMs using semantic data augmentation. SeaLLM focuses on the Southeast Asian (SEA) culture, which is adopted for Chinese and Korean cultures. TaiwanLLM focuses on traditional Chinese culture. Synatra-7B-v0.3-dpo and EEVE-Korean-10.8B-v1.0 targeted at Korean culture. CultureBank collects data from social media and we compare Arabic and Korean culture by fine-tuning GPT-3.5-turbo on its dataset.
@@
Main Results.
We analyzed the results from the culture and task sides in Figure 3. The most interesting observation is that our models outperformed GPT-4 on $5$ cultures and approached GPT-4 on the remaining $3$ cultures, although the data for fine-tuning are generated by GPT-3.5-turbo, which is much worse than GPT-4. We also generated cultural data via GPT-4 and fine-tuned other $8$ cultural-specific models for comparison, denoted as “Ours-gpt4” in Figure 3. The performance of those models is better than “Ours” (GPT-3.5-turbo version) but not so much. For other baselines, our models outperform them in most cases. Table 1 shows that our models achieved better performance than those costly LLMs which require pre-training and fine-tuning.
Ablation study.
Table 2 presents our ablation study on $4$ cultures, where “Generate” means just extracting opinions from the dialogue, “Verify” represents factually verifying the extracted opinions, and “Diversify” means removing redundant data. The results show that each module of CulturePark is effective, ensuring its interpretability.
@@ 4.2 Evaluation on cultural alignment via Hofstede’s Cultural Dimensions Theory
Results.
We compared our models (powered by GPT-3.5-Turbo) with GPT-3.5-turbo and GPT-4. As shown in Figure 4, our models outperform them by a large margin, indicating their excellent cultural alignment and cultural understanding abilities. Note that VSM is widely adopted as datasets for value and culture alignment, the results imply that our approach for data collection is effective, thus it could be further used for value alignment research.
@@ 4.3 Evaluation in situated learning for cultural education
Situated learning suggests that learning is best understood and facilitated when it occurs within the context (Anderson et al., 1996; Lave and Wenger, 1991). Motivated by situated learning, we leveraged CulturePark for cultural education where our fine-tuned models serve as foreigners to talk to people about cultural problems, which can create a situation for cross-cultural communication and learning cultural-specific knowledge. For example, a person who wants to learn about Arabic culture can communicate with our Arabic model.
Setup and study process.
We hired $24$ participants, each of whom was given an outline of cultural learning and asked to talk to models based on the outline. They can ask any related questions and express their opinions to models. Afterwards, the participants took a cultural understanding exam of VSM 2013 (Hofstede, 2013; Geert Hofstede, 2010) which they had never come into contact. We then computed the Euclidean distance between the ground truth and their answers from six cultural dimensions (rf. Section 4.2). For comparison, $12$ participants learned with our models and the other $12$ learned with GPT-4 to study $6$ cultures: Arabic, Bengali, German, Korean, Portuguese, and Spanish culture. Each culture was learned by four participants, two of them learning with our models and the others learning with GPT-4. Detailed information on the participants can be found in Table 9. During the study process, first, each participant was given an outline for cultural learning written by cultural experts. The outline (Section D.2), serves as the guideline for efficient learning. Then, we asked the participants to freely communicate with the models to learn about specific cultures. After the examination, we asked the participants to give a score of 1-5 to indicate their satisfaction with the learning process. In this study, our aim was to answer two questions: 1) What is the learning performance of the participants with our models and GPT-4? 2) How are their learning experience?
@@
Results.
Table 3 shows the averaged results of different participants. We have the following findings. First, participants learning with our models achieved better performance in cultural examination than those with GPT-4 in all cultures. This indicates that our fine-tuned models have a better cultural understanding than GPT-4. Second, participants are more satisfied with communicating with our models than GPT-4. Furthermore, many participants expressed that the responses from GPT-4 are vague. Even though we have prompted GPT-4 to be like a person from a specific culture, it always responds with neutral words that have no clear opinions or ideas. Instead, our models can provide straightforward opinions.
@@ 5.1 Why CulturePark benefits fine-tuning?
Cross-cultural communication vs. direct generation from GPT-4 / GPT-3.5. We compared the results of fine-tuning using data directly generated by GPT models (i.e., no communication). We generated such data by prompting GPT-4 or GPT-3.5-turbo as: “Question: {input} Answer: {output} Please list 10 reasons to support the answer and number them”. Then, these data are used to fine-tune GPT-3.5-turbo. LABEL:fig-directly shows the performance on content moderation tasks in Chinese, Korean and Turkish cultures. We see that data directly generated from GPT-4 is better than that from GPT-3.5, while our GPT-3.5-based models can outperform both of them.
Diversity of the generated data. We also analyzed the diversity gain (Bilmes, 2022) of the generated data for quality evaluation. We compared with CultureLLM (Li et al., 2024) and presented the results in LABEL:fig-diverse-gain. It indicates that CulturePark can generate more diverse and high-quality data.
@@ 5.2 Exploring agents’ cultural background and gender
To explore the influence of agent’s cultural background and gender, we conducted three types of multi-agent communications in Arabic culture: “In-cultural+Different gender”, “Cross-cultural+Same gender”, and “Cross-cultural+Different gender”. [^4] For a fair comparison, the number of seed data for this part is $50$ from WVS (Survey, 2022b) following CultureLLM (Li et al., 2024). For each setting, we fine-tuned three different models, whose training data is $500$, $750$, and $1000$, respectively. We evaluated the performance of the models on content moderation tasks and presented the results in LABEL:fig-gender-culture. “Cross-cultural+Different gender” exhibits the best performance and ability to generate more high-quality data. This indicates the necessity of bringing more diversity in data generation, as conducted in CulturePark.
@@
A potential dilemma arises when fine-tuning a large language model for specific tasks, as it may result in catastrophic forgetting of its original capabilities. This section explores the extent of forgetting exhibited by CulturePark on BIG-Bench-Hard (BBH) (Suzgun et al., 2022), which comprises 21 tasks that assess semantic understanding and logical reasoning. For cost efficiency, we sampled $100$ samples from each BBH task. We evaluated our models against the baseline model, GPT-3.5-turbo. The results in LABEL:fig-forget-BBH indicate that CulturePark generally maintains or even improves performance on most benchmarks, including the 21 tasks in BBH. This improvement suggests potential latent relationships between cultural data and general benchmarks, implying that fine-tuning of cultural data could improve general reasoning abilities.
@@ 5.4 Open-source fine-tuning with Llama2-70b
To verify the generalization ability of our framework, we leveraged the generated data to fine-tune cultural-specific Llama-2-70b models and evaluate on content moderation tasks. As shown in Figure 8, our models outperform Llama-2-70b in all $8$ cultures, especially in German, Chinese, Bengali and Portuguese cultures, which cover both low- and high-resource cultures. Furthermore, our models are also excellent in all $7$ tasks. [^5] The detailed hyperparameter setting is shown in Table 6. We only conduct this pilot study to show that Llama2 also works in CulturePark. The results might be further improved with more GPU resources. This verifies the generalization of CulturePark as an effective data collection platform.
@@ 6 Conclusions, Societal Impact, and Limitations
This paper introduced CulturePark, an LLM-powered multi-agent framework for cultural data collection through multi-agent communication. CulturePark can generate high-quality and diverse cross-cultural dialogue, which can be used to fine-tune culturally specific LLMs. We evaluated CulturePark across three downstream tasks: content moderation, cultural alignment, and cultural education, indicating great improvement over GPT-4.
@@
Our work has the following limitations. 1) More experiments can be performed by replacing GPT-3.5-Turbo in CulturePark to discover more results. 2) Our fine-tuned models are mostly for high-resource cultures. The reason is that the dataset and benchmark on low-resource cultures are rare, and we can not find enough data for fine-tuning and evaluation. 3) More efficient fine-tuning techniques can be studied to support the fine-tuning of culturally specific LLMs.
@@
The human study was conducted following local laws and regulations, and the evaluation process was controlled to ensure that no irresponsible content was generated. The authors respect all cultures studied in the world. The results of the paper may change due to the change in OpenAI API and its model versions.
Appendix A Discussion on the Relationship between Culture and Language
We strongly agree that language is not equal to, but only a part of culture. But using language to study culture is possible due to the following aspects:
  1. 1.
    Existing literature on culture understanding shows that culture boundaries are fluid, dynamic and uncertain. Delanoy emphasizes that cultures are not homogeneous or static entities but are fluid and dynamic. He critiques essentialist views that rigidly define cultural boundaries and instead promotes a more nuanced understanding that considers the intersections of various cultural factors, such as ethnicity, language, religion, and socio-economic conditions [Delanoy, 2020]. Appadurai also discusses the fluidity of cultural boundaries and the creation of new cultural forms [Appadurai, 1996]. Cultural boundaries can be geographical regions, language, religion and so on. Based on above statements, using language as cultural boundaries is reasonable.
  2. 2.
    Existing NLP works on culture also leverage labguage as culture boundaries. Naous et al. [2023] focuses on Arabic and English culture.  Wang et al. [2023] focuses on 8 different cultures: English, Chinese, French, Russian, German, Arabic, Japanese and Korean. Liu et al. [2023] also use language to split different cultures. The authors work on English, German, Russian, Bengali, Chinese, and Indonesian culture. Myung et al. [2024] is a hand-crafted benchmark for evaluate diverse cultures. They also use languages as culture boundaries.
  3. 3.
    Most downstream benchmarks are classified via language and we cannot get more fine-grained perspectives. For example, if we want to evaluate the performance of Arabic model, we can find benchmarks in Arabic culture. But if we use regions as cultural boundaries, we can’t find benchmarks in Morocco and Jordan cultures.
Appendix B Details on the cross-cultural dialogue
B.1 Examples on seed data
@@ B.2 The dialogue dataset
Then, we randomly sampled $750$ dialogues for each culture and evaluated the communication using prompts in Appendix E. As shown in Table 4, on average, the average ratio of statements that express cross-cultural understanding is $80.80$%. The analysis also verifies the effectiveness of CulturePark in extending topics and cross-cultural understanding.
B.3 Details on data refinement
Figure 11: Pipeline of data refinement.
Algorithm Figure 11 shows the algorithm pipeline of data refinement.
@@ B.3 Details on data refinement
Appendix C Details on the test sets
@@
Table 5: A brief introduction of the $8$ evaluation tasks and $41$ datasets. We list both the name and the size of test sets. For instance, “OSACT5(2541) [Mubarak et al., 2022]” denotes that there are 2541 test samples in the dataset OSACT5.
@@
Arabic (CulturePark-Ar) Middle East Offensive language detection: OSACT5(2541) [Mubarak et al., 2022]. Hate detection: Multi-Platform(675) [Chowdhury et al., 2020], OSACT5(2541) [Mubarak et al., 2022], and OSACT5_finegrained(2541) [Mubarak et al., 2022]. Vulgar detection: Multi-Platform(675) [Chowdhury et al., 2020] 8,973
@@
Bangli (CulturePark-Bn) Bangladesh Offensive language detection: TRAC2020 Task1(1000) [Bhattacharya et al., 2020], TRAC2020 Task2(1000) [Bhattacharya et al., 2020]. Hate detection: Hate Speech(1000) [Romim et al., 2021]. Threat detection: BACD(1000) [aimansnigdha, 2018]. Bias detection: BACD(1000) [aimansnigdha, 2018]. 5,000
@@
Chinese (CulturePark-Zh) China Spam detection: CCS(1000) [Jiang et al., 2019]. Bias detection: CDial-Bias(1000) [Zhou et al., 2022]. 2,000
@@
German (CulturePark-De) Germany and parts of Europe Offensive language detection: GermEval2018(3531) [Wiegand et al., 2018]. Hate detection: IWG_1(469) [Ross et al., 2016], IWG_2(469) [Ross et al., 2016], HASOC2020(850) [HASOC, 2020], and multilingual-hatecheck(1000) [Röttger et al., 2022]. 6,319
@@
Korean (CulturePark-Ko) South Korea Hate detection: hateSpeech(1000) [Moon et al., 2020], and HateSpeech2(1000) [daanVeer, 2020]. Abusive detection: AbuseEval(1000) [Caselli et al., 2020], CADD(1000) [Song et al., 2021], and Waseem(1000) [Waseem and Hovy, 2016]. 5,000
@@
Portuguese (CulturePark-Pt) Brazil and parts of Latin America Offensive language detection: OffComBR(1250) [de Pelle and Moreira, 2017], and HateBR(1000) [Vargas et al., 2022]. Bias detection: ToLD-Br-homophobia(1000) [Leite et al., 2020], and ToLD-Br-misogyny(1000) [Leite et al., 2020]. Abusive detection: ToLD-Br-insult(1000) [Leite et al., 2020]. 5,250
@@
Spanish (CulturePark-Es) Argentina, Mexico, and parts of Latin America Offensive language detection: AMI(1000) [Fersini et al., 2018], MEX-A3T(1000) [Álvarez-Carmona et al., 2018]. Hate detection: HatEval 2019(1000) [Basile et al., 2019]. Bias detection: DETOXIS_stereotype(1000) [de Paula and Schlicht, 2021], and DETOXIS_improper(1000) [de Paula and Schlicht, 2021]. Abusive detection: DETOXIS_abusive(1000) [de Paula and Schlicht, 2021], DETOXIS_mockery(1000) [de Paula and Schlicht, 2021]. Aggressive detection: DETOXIS_aggressiveness(1000) [de Paula and Schlicht, 2021]. Stance detection: DETOXIS_stance(1000) [de Paula and Schlicht, 2021]. 9,000
@@
Turkish (CulturePark-Tr) Turkey Offensive language detection: SemEval-2020(3528) [Zampieri et al., 2020], and offenseKaggle_2(1000) [Kaggle, 2022]. Abusive detection: ATC(1000) [Karayiğit et al., 2021]. Spam detection: Turkish Spam(825) [TurkishSpamV, 01]. Fine-grained offensive detection: offenseCorpus(1000) [Çöltekin, 2020]. 7,353
C.1 Arabic